OpenStudy (priyar):

Let f:R->[0,∞] be such that...the full q is below..

OpenStudy (priyar):

\[\lim_{x \rightarrow 5} ([f(x)]^2-9)/\sqrt{|x-5|}=0\] then find \[\lim_{x \rightarrow 5}f(x)\]

myininaya (myininaya):

hint bottom goes to 0 as x goes to 5 so top should go to 0 as x goes to 5

OpenStudy (priyar):

ya even i used the same so that we can apply lhospitals rule..but then i got f(x)=3..but thats not the correct answer..thats y i asked this q....

OpenStudy (freckles):

it could be \[\lim_{x \rightarrow 5} f(x)=3 \text{ or } \lim_{x \rightarrow 5}f(x)=-3 \] since f:R->[0,infty) then it should be 3

OpenStudy (priyar):

ya thats how i got 3..

OpenStudy (freckles):

oh wait

OpenStudy (priyar):

but was my reasoning as to why we should equate the numerator to zero correct @freckles ?

OpenStudy (freckles):

I think we need to go a step further

OpenStudy (freckles):

the bottom goes to 0

OpenStudy (freckles):

if you had 5/0 or -7/0 then the limit wouldn't exist to begin with

OpenStudy (freckles):

but if you have 0/0 then there is still a chance the limit will exist

OpenStudy (freckles):

we have taken into consideration only when the limit will exist now but we need to make sure the limit actually comes out to be 0

OpenStudy (priyar):

ya thats exactly what i thought too!

OpenStudy (priyar):

how can we do that?

OpenStudy (freckles):

(hmmm... only way for this limit to exist is if that one thing approaches 3) ...but... let's try to apply l'hospital I guess

OpenStudy (freckles):

hmmm but we know nothing about f'

OpenStudy (freckles):

are you sure the question is asking to find the limit as x goes to 5 of f(x)?

OpenStudy (freckles):

not f'(x) right?

OpenStudy (priyar):

ya sure..

OpenStudy (priyar):

but now i have a feeling the answer in the book might be wrong..

OpenStudy (freckles):

let's pretend they are asking for f' instead of f and let's see if we get the same answer in the back of the book...

OpenStudy (priyar):

why? if f(x ) is a const f'(x ) is zero right?

OpenStudy (freckles):

we do not know if f(x) is a constant we only know the limit of f(x) as x approaches 5

OpenStudy (priyar):

oh u r right

OpenStudy (priyar):

??

OpenStudy (freckles):

\[\frac{d}{dx} \sqrt{|x-5|}=\frac{d}{dx} \sqrt{x-5} \text{ assuming } x-5 > 0 \\ \frac{d}{dx} \sqrt{|x-5|}=\frac{1}{ 2 \sqrt{x-5}} ,x > 5 \\ \frac{d}{dx} \sqrt{|x-5|}=\frac{d}{dx}\sqrt{-(x-5)} \text{ assuming } x-5<0 \\ \frac{d}{dx} \sqrt{|x-5|}=\frac{-1}{2 \sqrt{-(x-5)}} \text{assuming } x<5 \\ \\ \text{ assume } x>5 \text{ then we have: } \\ \lim_{x \rightarrow 5^+} \frac{2 f'(x) f(x)}{\frac{1}{2 \sqrt{x-5}}} =\lim_{x \rightarrow 5^+} 2 f'(x) f(x) \cdot 2 \sqrt{x-5}\] and this approaches 0 no matter what because of the sqrt(x-5) thing so we wouldn't have enough info to find the limit as f' as x approaches 5 I don't think

OpenStudy (freckles):

what does the back of the book say the answer is?

OpenStudy (priyar):

2!!

OpenStudy (priyar):

i think its a printing mistake...i checked google but was not able to find the q to check the answer....

OpenStudy (freckles):

here is an example function f(x)=x-2 as x approaches 5 f(x) approaches 3 right?

OpenStudy (priyar):

ya

OpenStudy (priyar):

i have one more q that i need help in..i got some answer but it was wrong...can u help ?can i close this and post that?

OpenStudy (freckles):

\[\lim_{x \rightarrow 5}\frac{(x-2)^2-9}{\sqrt{|x-5|}}= \lim_{x \rightarrow 5} \frac{x^2-4x+4-9}{\sqrt{|x-5|}} \\ =\lim_{x \rightarrow 5} \frac{x^2-4x-5}{\sqrt{|x-5|}} \\ =\lim_{x \rightarrow 5} \frac{(x+1)(x-5)}{\sqrt{|x-5|}} \\ \text{ assume } x>5 \text{ then } |x-5|=x-5 \\ \text{ so we have } \\ \lim_{x \rightarrow 5^+} \frac{(x+1)(x-5)}{\sqrt{x-5}}=\lim_{x \rightarrow 5^+} (x-1) (x-5)^\frac{1}{2}\] and we see this goes to 0 as x goes to 5 from the right similarly we can find the left limit and we will also see this is 0

OpenStudy (freckles):

and again that was just me choosing a random function that went to 3 just to show the limit will be 0 with that condition and I agree your book seems to be wrong and okay next question

OpenStudy (freckles):

oops I change x+1 to x-1 but whatever

OpenStudy (priyar):

thank u